The one-sample t statistic from a sample of n = 13 observations for the two-sided test of the following hypotheses has the vaYou read that a study is planned for which a test of hypotheses will be done at significance level a size, the power is 0.95.
Answer #1

(a) The degrees of freedom fort is n-1 = 13 - 1 = 12. (b) At 10% level of significance and for 12 degrees of freedom, the two